OpenStudy (anonymous):

Should an estimate or the exact answer be calculated? Then do the appropriate calculation. For a science experiment, Renny had a container with 150 mL of solution. He transferred 24.8 mL to each of 3 smaller containers and 17.6 mL to each of 2 other small containers. How much solution was left in the original large container? A.The exact answer is needed. There is 40.4 mL of solution left in the large container. B. An estimate is sufficient. There is about 39 mL left in the large container.

OpenStudy (anonymous):

please help me

OpenStudy (tkhunny):

What's the trick? Exact Answer 150 - 24.8 - 24.8 - 24.8 - 17.6 - 17.6 = 150 - 3(24.8) - 2(17.6) = 150 - 74.4 - 35.2 = 40.4 One Possible Estimate 24.8 is about 25 17.6 is about 18 150 - 3(25) - 2(18) = 150 - 76 - 36 = 38 Another Possible Estimate 3(24.8) = 74.4 ==> 74 2(17.6) = 35.2 ==> 35 150 - 74 - 35 = 41 How else can you round?